OpenStudy (anonymous):

What % Of 6y is 24 Please show work i really need help :(

OpenStudy (anonymous):

Is this how the question is stated on your paper? seems a bit vague?

OpenStudy (anonymous):

Thank you for helping me Mathjerky :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not a problem. Just have two unknowns so it would be difficult to solve without another equation.

OpenStudy (anonymous):

is it 6y? or another number instead of the y?

OpenStudy (anonymous):

nope its just 6y

OpenStudy (anonymous):

cause that y means you can plug in any number there, so the possibilities of the % could be endless.

OpenStudy (anonymous):

They sid we could use a proportion to anwser the question in the terms of y

OpenStudy (anonymous):

General formula for this type is: is / of = % / 100, I believe. so, since 'is' is by the 24, is = 24 and 64 is next to the 'of', of = 24

OpenStudy (anonymous):

24 / 6y = % / 100

OpenStudy (anonymous):

just have two unknowns which is causing an issue. y & the %.

OpenStudy (anonymous):

did u use it as a proportion already?

OpenStudy (anonymous):

im just gunna assume they want it as a formula so you can just add in any 'y' value in the future to easily solve.

OpenStudy (anonymous):

so using 24 / 6y = % / 100, i would solve for the % to get the equation to find the %.

OpenStudy (anonymous):

you would want to get the % by itself, so multiply both sides by 100.

OpenStudy (anonymous):

then you get 100 * 24 / 6y = %

OpenStudy (anonymous):

you can then just write it as 2400 / 6y = %

OpenStudy (anonymous):

since 24 * 100 is 2400

OpenStudy (anonymous):

i think that is what they want for the answer. Because then for any value of 'y' you plug in, it w ill give you the %.

OpenStudy (anonymous):

What about What % of 7y is 119

OpenStudy (anonymous):

for example: y = 1, 2400 / 6(1) = % 2400 / 6 = % 400 = %

OpenStudy (anonymous):

this you can also use a proportion

OpenStudy (anonymous):

yeah, the same way, the other is set up.

OpenStudy (anonymous):

the formula that you use to plug in is: is / of = % / 100

OpenStudy (anonymous):

What # is 48% of 175

OpenStudy (anonymous):

same formula: is / of = % / 100 except this time they give you the % instead of the "is" so it would look like: is / 175 = 48 / 100

OpenStudy (anonymous):

then, you could just multiply 175 to both side to get "is" by itself.

OpenStudy (anonymous):

is = 175 * 48 / 100

OpenStudy (anonymous):

now this one confused me 150 is 15% of what #

OpenStudy (anonymous):

is / of = % / 100, is = 150, 15 = %. this one is just missing the "of" therefore, it would look like: 150 / of = 15 / 100

OpenStudy (anonymous):

however, in this case. it requires a slight bit more algebra to solve for the of. first, multiply of to get it to the numerator. 150 = of * 15 / 100 the you need to move the 15 / 100 to the other side, to do this we have to divide by 15 and multiple by 100 to cancel them on the right side.

OpenStudy (anonymous):

150 / 15 * 100 = of

OpenStudy (anonymous):

150 / 15 = 30, 30 * 100 = 3000

OpenStudy (anonymous):

oops, typo. its suppose to be 10, not 30.

OpenStudy (anonymous):

1000 is the answer.

OpenStudy (anonymous):

you always want to check your answers when you done, to make sure they make sense.

OpenStudy (anonymous):

you can do this by saying 15% of 1000 is what? 15% = .15 when you convert to decimals and 1000 * .15 = 150. which is what you started with.

OpenStudy (anonymous):

Hopefully, this helped. Gotta go help my friend! Good Luck!